OpenStudy (anonymous):

How do I solve this? (3.4 x 10^6)(6 x 10^3) / (5 x 10^5) How exactly do I simplify the numerator? Do I just add them? Multiply?

OpenStudy (anonymous):

\[\frac{ (3.4 x 10^6)(6 x 10^3)}{ (5 x 10^5) }\]

OpenStudy (skullpatrol):

$$\frac{(3.4 * 10^6)(6 * 10^3)}{(5 * 10^5)}= \frac{(3.4*10*10*10*10*10*10)*(6*10*10*10)}{5*10*10*10*10*10}$$

OpenStudy (skullpatrol):

By definition $$b^n=b*b*b*...*b$$n factors of b

OpenStudy (anonymous):

(3,400,000 x 6000)/(500,000)

OpenStudy (skullpatrol):

Yes, that is one way to do it, why not keep all the powers of 10 together?

OpenStudy (anonymous):

Uh. I'm not sure if I add or multiply when I keep the powers or 10 together. Is it (9.4 * 10^9)/(5∗10^5) or...

OpenStudy (skullpatrol):

Yes, you are correct there are 9 ten's being multiplied in the numerator and 6 times 3.4, right?

OpenStudy (skullpatrol):

$$6*3.4*10^9$$

OpenStudy (anonymous):

20.4 * 10^9

OpenStudy (skullpatrol):

Correct $$20.4*10^9$$is how you simplify the numerator.